OpenStudy (anonymous):

What are the intercepts of the line x- 3y = 9? x - intercept: - 3 ; y - intercept -9 x- intercept: 3 ; y - intercept 9 x- intercept: -9 ; y - intercept ; -3 or x- intercept: 9, y - intercept: -3

OpenStudy (anonymous):

answer ASAP!

OpenStudy (anonymous):

first put the equation in the form of y=mx+c

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

0 - 3y = 9 is - 3 and x - 3 0= 9 is 9 thats what i got

OpenStudy (anonymous):

yea right :)

OpenStudy (anonymous):

okay thanks you two :)